Kritika Roy

816 East Kerr Avenue, Apt 205  Cell: (217) 419-1243

Urbana - 61802 IL, USA E-mail: drkritika.roy@gmail.com Contact Information

Objective To obtain a job that will give me an insight into the work, matches my interests and enhances my skills.

Education University-Rajasthan: University of Health Science, Jaipur, Rajasthan India B.D.S. (Bachelor in Dental Surgery) Jaipur Dental College, 2004 - 2009

• First Year : 2004-05, First Attempt, 74.83%

• Second Year: 2005-06, First Attempt, 82.33%

• Third Year : 2006-07, First Attempt, 74.62%

• Fourth Year: 2007-08, First Attempt, 72.64%

• Internship : 2009

Professional Bharadwaj Dental Clinic, Jaipur India Experience

Dental Assistant and Clinic Manager August 2009 to August 2011

• Worked as a Dental Assistant at a Dental Clinic

• Managed the Administrative tasks at the Clinic Self Dental Practice

Dental Practice and Clinic Manager August 2011 to March 2013

• Part-time Dental Practice in India

Awards and Achievements • Among the top ten University Toppers for consecutive years during BDS for the

session 2004-05, session 2005-06 and the session 2006-07.

• Presented a paper on Dental Hygiene at a conferece in India during the session 2006-07.

• Was Prefect of School for the year 2003-04.

Technical Skills Computer Applications: most common productivity packages like Office, Internet.

Age: 27 years Personal Details

Work Permit: Yes Interests: I love helping people, especially children and like playing with and teaching small children, I also like flower decoration and home decoration and health consciousness

1 of 1